Consider the sequence 130, 143, 156, 169, ...

Write an explicit formula to represent the arithmetic sequence and use it to find the 13th term.


A. A(n) = 130 + (n-1)13; 286
B. A(n) = 130 + 13n; 299
C. A(n) = 130 + (n-1)13; 299
D. A(n) = 130 + 13n; 286

Answer by richwmiller(17219)

A. A(n) = 130 + (n-1)13; 286
A(13) = 130 + (12)13
A(13) = 130 + 156
A(13) = 286